A police jeep starts chasing a culprit with speed of 90km/h and acceleration 2m/sculprit is moving with 108 km/h culprit has already moved 300m extra find when and where they will meet
 

Grade:12th pass

1 Answers

Himanshu Rathour
5 years ago
Relative velocity of jeep wrt culprit=v =108-90=18km/h
=5m/s
Relative acceleration =-2m/s
Relative distance =-300
Let they meet after time t
-300=5t-t²
t²-5t-300=0
t=20s
Distance covered by culprit in 20 sec=20*25
=500m
Hence they meet after 800m from point of start of jeep